OpenStudy (anonymous):

Add v+t/vt^2+6v+t/v^2t

OpenStudy (anonymous):

\[\frac{v+t}{vt^2}+\frac{6v+t}{v^2t}\] \[\frac{v(v+t)+t(6v+t)}{v^2t^2}\] \[\frac{v^2+vt+6vt+t^2}{v^2t^2}\] \[\frac{v^2+7vt+t^2}{v^2t^2}\]